# XLD: A Cross-Lane Dataset for Benchmarking Novel Driving View Synthesis

### v1

@@ Abstract
Thoroughly testing autonomy systems is crucial in the pursuit of safe autonomous driving vehicles. It necessitates creating safety-critical scenarios that go beyond what can be safely collected from real-world data, as many of these scenarios occur infrequently on public roads. However, the evaluation of most existing NVS methods relies on sporadic sampling of image frames from the training data, comparing the rendered images with ground truth images using metrics. Unfortunately, this evaluation protocol falls short of meeting the actual requirements in closed-loop simulations. Specifically, the true application demands the capability to render novel views that extend beyond the original trajectory (such as cross-lane views), which are challenging to capture in the real world. To address this, this paper presents a novel driving view synthesis dataset and benchmark specifically designed for autonomous driving simulations. This dataset is unique as it includes testing images captured by deviating from the training trajectory by $1-4$ meters. It comprises six sequences encompassing various time and weather conditions. Each sequence contains $450$ training images, $150$ testing images, and their corresponding camera poses and intrinsic parameters. Leveraging this novel dataset, we establish the first realistic benchmark for evaluating existing NVS approaches under front-only and multi-camera settings. The experimental findings underscore the significant gap that exists in current approaches, revealing their inadequate ability to fulfill the demanding prerequisites of cross-lane or closed-loop simulation. Our dataset is released publicly at the project page: https://3d-aigc.github.io/XLD/.
@@ 1 Introduction
In closed-loop simulation, it is essential for the vehicle to have the freedom to respond to control commands within the simulation environment, rather than being confined to strictly following the original trajectory from logs. To corporate with the recently developed end-to-end autonomous driving [55, 21], designing a neural simulator for AD simulation [52, 48], which can render photo-realistic images on novel views for closed-loop simulation and algorithm training, is in high demands. The main scientific problems boil down to the 3D reconstruction [28] and novel view synthesis (NVS) [50, 48], which are also long-standing problems in computer vision and computer graphics. Traditional methods such as [35, 36] have dominated the major deployment of 3D scene reconstruction for a long time. However, those reconstructed scenes cannot be directly used to produce photo-realistic novel views, thus imposing large restrictions on sensor simulations. As a result, the industrial bridges the sim-to-real gap by parametric and procedural modeling technique [42] or human-involved creations. With the recent rapid development of 3D implicit field, such as neural radiance field (NeRF [30]) and explicit primitive representations, i.e., 3D Gaussian Splatting (3DGS [23]), reconstructing a scene from a collection of images serves as the foundation of end-to-end autonomous driving simulation [20]. These techniques enable rendering high-quality and photorealistic images on novel views.
@@ 2.2 NeRF-based NVS for Driving Simulation
Motivated by the NeRF-based methods and point-based differentiable rendering [53, 1, 34], 3D Gaussian Splatting (3DGS) [23] opens a new era with the leading advantages in explicit representation and real-time rendering capability. Within a concise timeframe, numerous methods have emerged that focus on road scene reconstruction and NVS by leveraging the 3DGS representation. For instance, PVG [12] introduces periodic vibration-based temporal dynamics to reconstruct dynamic urban scenes. StreetGaussian [50] models the dynamic urban street environment as a collection of point clouds with semantic logits and 3D Gaussians, each associated with either a foreground vehicle or the background. DrivingGaussian [56] uses incremental static 3D Gaussians to represent the scene’s static background. It also employs a composite dynamic Gaussian graph to handle multiple moving objects with LiDAR data.  [26] introduce a hybrid method that combines radiance fields with 3DGS representation, effectively eliminating the requirement for point initialization in the context of urban scene NVS.  [47] propose a hybrid Gaussian representation specifically designed for performing online dense mapping in unbounded large-scale scenes. GaussianPro [13] utilizes priors from reconstructed scene geometries and patch matching techniques to generate precise Gaussians, leveraging the scene’s existing structure. DC-Gaussian [45] introduces adaptive image decomposition for modeling reflections and occlusions. It incorporates illumination-aware obstruction modeling to handle reflections and occlusions under varying lighting conditions in urban scene novel view synthesis. Our dataset and benchmark specifically focus on evaluating the performance of neural-based driving simulation in NVS, particularly in cross-lane scenarios. A portion of the mentioned methods serve as our baselines, taking into account the code availability.
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
The authors would like to thank Yuanyuan Gao and Jingfeng Li for preparing the experiments. This work was funded by Baidu Inc.

### v2

@@ Abstract
Thoroughly testing autonomy systems is crucial in the pursuit of safe autonomous driving vehicles. It necessitates creating safety-critical scenarios that go beyond what can be safely collected from real-world data, as many of these scenarios occur infrequently on public roads. However, the evaluation of most existing NVS methods relies on sporadic sampling of image frames from the training data, comparing the rendered images with ground truth images using metrics. Unfortunately, this evaluation protocol falls short of meeting the actual requirements in closed-loop simulations. Specifically, the true application demands the capability to render novel views that extend beyond the original trajectory (such as cross-lane views), which are challenging to capture in the real world. To address this, this paper presents a novel driving view synthesis dataset and benchmark specifically designed for autonomous driving simulations. This dataset is unique as it includes testing images captured by deviating from the training trajectory by $1-4$ meters. It comprises six sequences encompassing various time and weather conditions. Each sequence contains $450$ training images, $120$ testing images, and their corresponding camera poses and intrinsic parameters. Leveraging this novel dataset, we establish the first realistic benchmark for evaluating existing NVS approaches under front-only and multi-camera settings. The experimental findings underscore the significant gap that exists in current approaches, revealing their inadequate ability to fulfill the demanding prerequisites of cross-lane or closed-loop simulation. Our dataset is released publicly at the project page: https://3d-aigc.github.io/XLD/.
@@ 1 Introduction
In closed-loop simulation, it is essential for the vehicle to have the freedom to respond to control commands within the simulation environment, rather than being confined to strictly following the original trajectory from logs. To corporate with the recently developed end-to-end autonomous driving [61, 23], designing a neural simulator for AD simulation [58, 54, 52], which can render photo-realistic images on novel views for closed-loop simulation and algorithm training, is in high demands. The main scientific problems boil down to the 3D reconstruction [33] and novel view synthesis (NVS) [56, 54], which are also long-standing problems in computer vision and computer graphics. Traditional methods such as [40, 41] have dominated the major deployment of 3D scene reconstruction for a long time. However, those reconstructed scenes cannot be directly used to produce photo-realistic novel views, thus imposing large restrictions on sensor simulations. As a result, the industrial bridges the sim-to-real gap by parametric and procedural modeling technique [47] or human-involved creations. With the recent rapid development of 3D implicit field, such as neural radiance field (NeRF [35]) and explicit primitive representations, i.e., 3D Gaussian Splatting (3DGS [25]), reconstructing a scene from a collection of images serves as the foundation of end-to-end autonomous driving simulation [22]. These techniques enable rendering high-quality and photorealistic images on novel views.
@@ 2.2 NeRF-based NVS for Driving Simulation
Motivated by the NeRF-based methods and point-based differentiable rendering [59, 1, 39, 30], 3D Gaussian Splatting (3DGS) [25] opens a new era with the leading advantages in explicit representation and real-time rendering capability. Within a concise timeframe, numerous methods [14, 56] have emerged that focus on road scene reconstruction and NVS by leveraging the 3DGS representation. For instance, PVG [14, 27] introduces periodic vibration-based temporal dynamics to reconstruct dynamic urban scenes. StreetGaussian [56] models the dynamic urban street environment as a collection of point clouds with semantic logits and 3D Gaussians, each associated with either a foreground vehicle or the background. DrivingGaussian [62] uses incremental static 3D Gaussians to represent the scene’s static background. It also employs a composite dynamic Gaussian graph to handle multiple moving objects with LiDAR data.  [31] introduce a hybrid method that combines radiance fields with 3DGS representation, effectively eliminating the requirement for point initialization in the context of urban scene NVS.  [53] propose a hybrid Gaussian representation specifically designed for performing online dense mapping in unbounded large-scale scenes. GaussianPro [15] utilizes priors from reconstructed scene geometries and patch-matching techniques to generate precise Gaussians, leveraging the scene’s existing structure. DC-Gaussian [50] introduces adaptive image decomposition for modeling reflections and occlusions. It incorporates illumination-aware obstruction modeling to handle reflections and occlusions under varying lighting conditions in urban scene novel view synthesis. Our dataset and benchmark specifically focus on evaluating the performance of neural-based driving simulation in NVS, particularly in cross-lane scenarios. A portion of the mentioned methods serve as our baselines, taking into account the code availability. Moreover, generalizable 3D-GS methods such as PixelSplat [11], Mvsplat [13], and GGRt [26] also attempts to synthesis novel images within a well-pretrained generalizable feed-forward Gaussian networks.
@@ 5 Conclusion
Acknowledgments
The authors would like to thank Yuanyuan Gao and Jingfeng Li for preparing the experiments.